MatthewCompany purchased a new machine on October 1, 2017, at a cost of $145,000. The company estimated that the machine will have a salvage value of $25,000. The machine is expected to be used during its 5-year life.

Instructions

Compute the depreciation expense under the following methods for the year indicated. (a) Straight-line for 2017. (b) Declining-balance using double the straight-line rate for 2017 and 2018.
Answer #1

Requirement a

Depreciation for 2017= $ 6,000

.

Working

Straight line Method
A Cost $ 145,000
B Residual Value $ 25,000
C=A - B Depreciable base $ 120,000
D Life [in years left ]                                  5
E=C/D Annual SLM depreciation $             24,000.00

.

Depreciation schedule-Straight line method
Year Book Value Depreciation expense Accumulated Depreciation Ending Book Value
2017 $        145,000.00 $               6,000.00* $                6,000.00 $            139,000.00
2018 $        139,000.00 $             24,000.00 $             30,000.00 $            115,000.00

*For 3 months.

Requirement b

Depreciation for 2017= $ 14,500

Depreciation for 2018= $ 52,200

Working

Double declining Method
A Cost $ 145,000
B Residual Value $ 25,000
C=A - B Depreciable base $ 120,000
D Life [in years] 5
E=C/D Annual SLM depreciation $ 24,000
F=E/C SLM Rate 20.00%
G=F x 2 DDB Rate 40.00%

.

Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
2017 $ 145,000 40.00% $ 14,500 $ 14,500 $ 130,500
2018 $ 130,500 40.00% $ 52,200 $ 66,700 $ 78,300